# DextrAH-G: Pixels-to-Action Dexterous Arm-Hand Grasping with Geometric Fabrics

### v2

@@ F.1 Depth Image Details
We simulation-rendered depth images are entirely free of noise as shown in Figure 11. To address the visual sim2real gap, we add a variety of augmentations to the rendered depth image. Specifically, the following augmentations are added: 1) a pixel value would be set to 0 with probability $p_{dropout}=0.003$, 2) a pixel value would be set to a random value $\in(-0.5,-1.3)$ with probability $p_{randu}=0.003$, 3) a linear segment of pixels of up to 18 pixels length and 3 pixels width would appear with probability of $p_{stick}=0.0025$ to mimic robot wires and other artifacts, and 4) the uncorrelated and correlated depth noise models and their parameters are used exactly as reported in [57]. The effect of these augmentations can be seen in Figure 11. See our video to better visualize the depth image noise in both simulation and the real-world.
@@ F.4 Simulation Experiments
We evaluate our teacher and student policy in the simulator across 140 different objects from the training dataset. The objects are visualized in Figure 6. The policy has 10 seconds window to grasp the object. If the object is lifted to the target position in the air, the attempt is labeled as successful. Otherwise, it’s a failed attempt. Results are shown in Figure 12.
@@ Appendix G DextrAH-G State Machine for Bin Packing
Appendix H Additional Real Experiment Analysis
@@
A time-series plot of CS is shown in Figure  14. Overall, DextrAH-G consistently grasps and transports several objects in a row before experiencing a failure. Several times, DextrAH-G moved more than 20 objects in a row before experiencing a failure. DextrAH-G did produce failures within 5 CS several times due to the previously mentioned failures modes.
Figures  15 and  16 report DextrAH-G’s speed performance index. Overall, DextrAH-G successfully grasps and transports 5.64 objects per minute and DextrAH-G handled about 42% of objects within 6 s - 7.3 s. With persistent uptime, dexterous grasping with this speed level becomes useful for real-world applications.
@@
Finally, we scope DextrAH-G’s signals over a short horizon during grasping and plot the results in Figure 18. As seen, the depth FGP can output quickly shifting actions while the geometric fabric produces smooth joint angle targets. Moreover, since the desired velocity is set to zero for the PD controller, the arm’s physical position lags behind the targets about 0.2 s. The hand’s physical position lags behind the targets by about 0.1 s. Despite this phase lag, DextrAH-G operates well due to RL training for these lagged dynamics. We also see increased separation between desired and measured joint angle positions during the later phase of DextrAH-G execution because the robot is grasping the object. The increased tracking error induces contact forces between the robot and the object which facilitates grasping. Overall, the hand experiences greater joint angle swings during execution than the arm. This quality is desirable and indicates that the hand is doing much of the mechanical work required for grasping while minimally moving the arm. It is also the case that arm joint angle movements translate to much greater Cartesian palm movement so only minimal arm angular movement is required for effectively guiding the hand during grasping.
@@
Figure 15: From aggregating 256 attempted grasp-and-transports across 30 objects, we demonstrate 5.64 successful grasp-and-transports per minute, with 0.53 interventions per minute and 0.30 failures per minute. We define successful grasp-and-transport as the robot taking actions that result in the object being grasped and lifted off the table, and then dropped into the bin. We define interventions as the human adjusting the position or orientation of the object after the robot had failed to grasp the object. We define failures as the robot taking actions that result in the object falling off the table or moving far away enough that the human decided to move onto another object instead of intervening. If the robot misses a grasp or drops the object, but then subsequently retries successfully without human intervention, this counts as one successful grasp-and-transport, zero interventions, and zero failures. Note that these rates are aggregate results of total cumulative count divided by total time, so this rate of successful grasp-and-transports is affected by the time taken for interventions and failures.
@@
Figure 17: We visualize real-world per-object results over continuous runs across 30 objects. DextrAH-G performed best on real-world objects such as the brick, spam, duck, mouse box, and container. It performed least well on real-world objects such as the small bottle, pot, and green cup. We define successful grasp-and-transport as the robot taking actions that result in the object being grasped and lifted off the table, and then dropped into the bin. We define intervention as the human adjusting the position or orientation of the object after the robot had failed to grasp the object. We define failure as the robot taking actions that result in the object falling off the table or moving far away enough that the human decided to move onto another object instead of intervening. If the robot misses a grasp or drops the object, but then subsequently retries successfully without human intervention, this counts as one successful grasp-and-transport, zero interventions, and zero failures.

### v3

@@ F.1 Depth Image Details
We simulation-rendered depth images are entirely free of noise as shown in Figure 13. To address the visual sim2real gap, we add a variety of augmentations to the rendered depth image. Specifically, the following augmentations are added: 1) a pixel value would be set to 0 with probability $p_{dropout}=0.003$, 2) a pixel value would be set to a random value $\in(-0.5,-1.3)$ with probability $p_{randu}=0.003$, 3) a linear segment of pixels of up to 18 pixels length and 3 pixels width would appear with probability of $p_{stick}=0.0025$ to mimic robot wires and other artifacts, and 4) the uncorrelated and correlated depth noise models and their parameters are used exactly as reported in [57]. We also stochastically perturbed the depth camera placement in simulation during distillation so that the distilled FGP gained robustness to calibration errors in the real world. The effect of these augmentations can be seen in Figure 13. See our video to better visualize the depth image noise in both simulation and the real-world.
@@ F.4 Simulation Experiments
We evaluate our teacher and student policy in the simulator across 140 different objects from the training dataset. The objects are visualized in Figure 8. The policy has a 10 second window to grasp the object. If the object is lifted to the target position in the air, the attempt is labeled as successful. Otherwise, it’s a failed attempt. Results are shown in Figure 14.
@@ Appendix G DextrAH-G State Machine for Bin Packing
Appendix H Nomenclature
We summarize our nomenclature in Table 3.
3.1 Geometric Fabrics and Fabric-Guided Policies (FGPs)
Symbol Meaning
$\mathbf{M}_{f}\in\mathbb{R}^{n\times n}$ Positive-definite system metric (mass), which captures system prioritization.
$\mathbf{f}_{f}\in\mathbb{R}^{n}$ Nominal path generating geometric force.
$\mathbf{f}_{\pi}(\mathbf{a})\in\mathbb{R}^{n}$ Additional driving force of some action $\mathbf{a}\in\mathbb{R}^{m}$.
$\mathbf{q}_{f},{\dot{\mathbf{q}}}_{f},{\ddot{\mathbf{q}}}_{f}\in\mathbb{R}^{n}$ Position, velocity, and acceleration of the fabric ($\mathbf{q}_{f}$ used as PD target to control robot).
$\mathbf{x}=\phi_{fk}(\mathbf{q})\in\mathbb{R}^{3}$ Origin of each sphere used to model the robot geometry (computed through forward kinematics).
$\mathbf{r}_{i}\in\mathbb{R}^{3}$ Closest point on collision body $i$ to a given robot body sphere.
$\hat{\mathbf{n}}_{i}\in\mathbb{R}^{3}$ Direction from a given robot body sphere to the closest point on collision body $i$.
$d_{i}\in\mathbb{R}$ Signed distance between a given robot body sphere and collision body $i$.
$\underline{d}_{i}\in\mathbb{R}^{+}$ Lower-bounded distance between a given robot body sphere and collision body $i$.
${\ddot{\mathbf{x}}}_{b}\in\mathbb{R}^{3}$ Base acceleration response per robot body sphere away from collision.
$\mathbf{M}_{b}\in\mathbb{R}^{3\times 3}$ Base metric response per robot body sphere for collision avoidance.
$s_{i}\in\mathbb{R}$ Smooth velocity gate that goes high when this robot body sphere is moving towards collision body $i$.
$v_{i}\in\mathbb{R}$ Signed impact speed that is negative when moving towards collision body $i$.
$\mathbf{A}\in\mathbb{R}^{5\times 16}$ First five components from PCA on grasping motion data retargeted from human hand to Allegro hand.
$\mathbf{x}_{f,target}\in\mathbb{R}^{3}$ Target palm position.
$\mathbf{r}_{f,target}\in\mathbb{R}^{3}$ Target palm orientation in Euler angles.
$\mathbf{x}_{pca,target}\in\mathbb{R}^{5}$ Target PCA position for the fingers.
3.2 Teacher Privileged FGP Training (Reinforcement Learning)
Symbol Meaning
$\mathbf{o}_{privileged}$ A subset of privileged state information provided to the teacher policy.
$\pi_{privileged}(\mathbf{o}_{privileged})$ Teacher policy that is trained with $\mathbf{o}_{privileged}$.
$\mathbf{s}$ All privileged state information provided to the critic.
$V(\mathbf{s})$ Critic value function that is trained with $\mathbf{s}$.
$\mathbf{o}_{robot}$ Robot state information.
$\mathbf{x}_{goal}\in\mathbb{R}^{3}$ Goal object position.
$\mathbf{o}_{obj}$ Object state information provided to the teacher policy.
$\widetilde{\mathbf{x}}_{obj}\in\mathbb{R}^{3}$ Noisy object position.
$\widetilde{\mathbf{q}}_{obj}\in\mathbb{R}^{4}$ Noisy object quaternion.
$\mathbf{e}\in\{0,1\}^{N_{objects}}$ Object one-hot embedding.
$\mathbf{x}_{palm},\mathbf{x}_{palm-x},\mathbf{x}_{palm-y}\in\mathbb{R}^{3}$ Positions of three points on the palm.
$\mathbf{x}_{fingertips}\in\mathbb{R}^{N_{fingers}\times 3}$ Positions of the fingertips.
$\mathbf{s}_{privileged}\in\mathbb{R}^{n}$ Privileged state information provided only to the critic.
$\mathbf{f}_{dof}\in\mathbb{R}^{N_{q}}$ Robot joint forces.
$\mathbf{f}_{fingers}\in\mathbb{R}^{N_{fingers}\times 3}$ Fingertip contact forces.
$\mathbf{x}_{obj}\in\mathbb{R}^{3}$ True object position.
$\mathbf{q}_{obj}\in\mathbb{R}^{4}$ True object quaternion.
$\mathbf{v}_{obj}\in\mathbb{R}^{3}$ True object velocity.
$\mathbf{w}_{obj}\in\mathbb{R}^{3}$ True angular velocity.
$\mathbf{a}\in\mathbb{R}^{11}$ Policy action, which is passed to the underlying geometric fabric.
$\mathcal{L_{PPO}}\in\mathbb{R}$ PPO loss.
3.3 Student Depth FGP Training (Policy Distillation)
Symbol Meaning
$\mathbf{o}_{depth}$ Observation provided to the student policy.
$\pi_{depth}(\mathbf{o}_{depth})$ Student policy that is trained with $\mathbf{o}_{depth}$.
$\mathbf{I}\in[0.5,1.5]^{160\times 120}$ m Raw depth image.
$\hat{\mathbf{a}}\in\mathbb{R}^{11}$ Student predicted actions.
$\hat{\mathbf{x}}_{obj}\in\mathbb{R}^{3}$ Predicted object position.
$\mathcal{L}\in\mathbb{R}$ Student supervision loss.
$\mathcal{L}_{action}\in\mathbb{R}$ Student action loss.
$\mathcal{L}_{pos}\in\mathbb{R}$ Student position loss.
$\beta\in\mathbb{R}$ Weight for student position loss.
Table 3: Nomenclature used in Section 3. Symbols used in Figure 2 are highlighted in yellow.
Appendix I Additional Real Experiment Analysis
@@
A time-series plot of CS is shown in Figure  16. Overall, DextrAH-G consistently grasps and transports several objects in a row before experiencing a failure. Several times, DextrAH-G moved more than 20 objects in a row before experiencing a failure. DextrAH-G did produce failures within 5 CS several times due to the previously mentioned failures modes.
Figures  17 and  18 report DextrAH-G’s speed performance index. Overall, DextrAH-G successfully grasps and transports 5.64 objects per minute and DextrAH-G handles 42% of objects within 6 s - 7.3 s. With persistent uptime, dexterous grasping with this speed level becomes useful for real-world applications.
@@
In tandem with the results in Figure 8, we find that the primary factors that affect policy performance include object size (small bottle), object slipperiness (green cup), object’s propensity to easily roll (apple), transparent objects (sanitizer bottle), and geometric aspects of the object that increase the likelihood of the fingers “catching” or “snagging” on the object.
Finally, we scope DextrAH-G’s signals over a short horizon during grasping and plot the results in Figure 20. As seen, the depth FGP can output quickly shifting actions while the geometric fabric produces smooth joint angle targets. Moreover, since the desired velocity is set to zero for the PD controller, the arm’s physical position lags behind the targets by about 0.2 s. The hand’s physical position lags behind the targets by about 0.1 s. Despite this phase lag, DextrAH-G operates well due to RL training for these lagged dynamics. We also see increased separation between desired and measured joint angle positions during the later phase of DextrAH-G execution because the robot is grasping the object. The increased tracking error induces contact forces between the robot and the object which facilitates grasping. Overall, the hand experiences greater joint angle swings during execution than the arm. This quality is desirable and indicates that the hand is doing much of the mechanical work required for grasping while minimally moving the arm. It is also the case that arm joint angle movements translate to much greater Cartesian palm movement so only minimal arm angular movement is required for effectively guiding the hand during grasping.
@@
Figure 17: From aggregating 256 attempted grasp-and-transports across 30 objects, we demonstrate 5.64 successful grasp-and-transports per minute, with 0.53 interventions per minute and 0.30 failures per minute. We define successful grasp-and-transport as the robot taking actions that result in the object being grasped and lifted off the table, and then dropped into the bin. We define interventions as the human adjusting the position or orientation of the object after the robot has failed to grasp the object. We define failures as the robot taking actions that result in the object falling off the table or moving far away enough that the human decides to move onto another object instead of intervening. If the robot misses a grasp or drops the object, but then subsequently retries successfully without human intervention, this counts as one successful grasp-and-transport, zero interventions, and zero failures. Note that these rates are aggregate results of total cumulative count divided by total time, so this rate of successful grasp-and-transports is affected by the time taken for interventions and failures.
@@
Figure 19: We visualize real-world per-object results over continuous runs across 30 objects. DextrAH-G performed best on real-world objects such as the brick, spam, duck, mouse box, and container. It performed least well on real-world objects such as the small bottle, pot, and green cup. We define successful grasp-and-transport as the robot taking actions that result in the object being grasped and lifted off the table, and then dropped into the bin. We define intervention as the human adjusting the position or orientation of the object after the robot has failed to grasp the object. We define failure as the robot taking actions that result in the object falling off the table or moving far away enough that the human decides to move onto another object instead of intervening. If the robot misses a grasp or drops the object, but then subsequently retries successfully without human intervention, this counts as one successful grasp-and-transport, zero interventions, and zero failures.